OpenStudy (anonymous):

Variables on each side. One hotel charges $1 plus $0.80 per minute for long distance and another hotel charges $2 plus $0.75 for long distance calls. Find the length of a call for which the two hotels charge the same amount.

OpenStudy (watchmath):

Let \(x\) be the length of a call. Then \(1+0.8x=2+0.75x\) \(0.05x=1\) \(x=1/0.05=20\) minutes